# FOSS: A Self-Learned Doctor for Query Optimizer

### v1

@@ VI-B3 Optimization Time Comparison
We compare the optimization time of various optimizers on the entire JOB benchmark, which is the time from the input of SQL to the generation of the execution plan. As depicted in Fig. 6, we represent the comparative results using box plots. In contrast to PostgreSQL, learned-based optimizers often require longer optimization time due to operations like model inference and plan encoding. However, the overhead is negligible compared to the time saved in execution. Among these learned-based optimizers, FOSS has the smallest 25th percentile, 50th percentile, and 75th percentile of optimization time, and its distribution is relatively concentrated. Most of its optimization time is spent on obtaining the original plan and plan encoding. Excessive reliance on the expert optimizer results in long optimization time for Bao and HybridQO. Additionally, Balsa, characterized by a lengthy average decision sequence and a wide range of plan exploration strategies, also performs poorly in terms of optimization time.
@@
VI-B4 Known Best Plan Comparison
In this subsection, we analyze the optimization upper bounds of each learned-based query optimizer. We conduct five runs for each method on the entire JOB workload. From all the results, we acquire the known best plan (i.e., the one with the lowest execution latency) on each query for each method. The comparison results are shown in Fig. 7.
It illustrates the time savings ratios of the learned-based methods relative to expert optimizer in terms of known best plans. Overall, FOSS comprehensively outperforms the other three methods, surpassing the expert optimizer on a broader range of queries at each percentage point. Specially, FOSS, HybridQO, Balsa, and Bao achieve at least a $25\%$ time savings on 78, 54, 45, and 35 queries, respectively. They each save at least $75\%$ of the time on 33, 30, 23, and 14 queries, respectively. Additionally, due to the limited plan search space of Bao, it generates better-performance plans than expert optimizer across a minimal range of queries. Balsa, lacking assurance from the original plan, generates plans with poorer-performance than expert optimizer on many queries.
@@ VI-C Analysis of design choices
2-Maxsteps 8.52 152.95 0.46
3-Maxsteps (FOSS) 9.23 170.57 0.44
4-Maxsteps 11.24 178.18 0.48
5-Maxsteps 15.21 190.43 0.53
Off-Simulated 48.01 171.12 0.65
Off-Validation 7.90 173.82 0.62
2-Agents 14.21 202.25 0.42
@@ VI-C1 Determination of Maxsteps
Larger $maxsteps$ implies that FOSS can explore more diverse plans through the modifications of longer sequences to the original plan. However, it also makes the training of FOSS more challenging. We set $maxsteps$ to 2, 3, 4, and 5, respectively, and compare their performances. As described in Table II, with an increase in $maxsteps$, the optimization time grows due to the consideration of more plans. Concurrently, training time gradually increases as the plan search space expands, and the number of plans to be executed grows. As illustrated in Fig. 8, when $maxsteps$ is set to 3, FOSS achieves the best performance with the minimum GMRL.
For a more in-depth comparison, as described in Fig. 6, we analyze the distribution of step status for known best plan, where step0 plans refers to the original plans, and stepN plans refers to plans that take $N$ steps from the original plans. In the bar for $2-maxsteps$, there is a significant number of step2 plans. It suggests that $2-maxsteps$ may be insufficient. In the bar for $5-maxsteps$, the proportion of step4 and step5 plans is relatively small. Despite spending nearly twice the training time of $2-maxsteps$, $5-maxsteps$ yields the lowest GMRL. This is because as $maxsteps$ increases, the longer action sequences complicates the agent’s exploration and makes it more challenging for the $\theta_{adv}$ to select a good plan from candidate plans. As demonstrated by the other two results, FOSS always yields effective plans within 1-3 steps. Considering both training complexity and experimental effectiveness, we set the $maxsteps$ to 3.
@@ VI-C2 Effect of Simulated Environment
To illustrate the effect of the simulated environment, we disable it in FOSS. It allows the agent to interact only with the real environment. Due to the extensive queries executed in the real environment, it significantly prolong the training time. To make training more feasible, we have to reduce the query number for each agent update to 200. After 48 hours of training, despite benefiting from more reliable rewards feedback, its performance remains unsatisfactory. Its final GMRL is 0.63, which is significantly poorer than FOSS. The inadequacy can be ascribed to the insufficient exploration. It indicates the effectiveness of the simulated environment.
VI-C3 Effect of Validation Phase
We deploy FOSS without the validation of estimated optimal plans. As described in Table II, it reduces training time but diminishes the performance. As shown in Fig. 8, its GMRL decreases slowly with training progress, even performing worse than the expert optimizer in the early stages of training. The absence of the validation phase reduces diversity in $\theta_{adv}$ training data and prevents timely error alleviation, resulting in the accumulation of errors during DRL training.
VI-C4 Training with Muti-Agents
To enhance the robustness of FOSS, we also incorporate a switch for multi-agents. FOSS can initialize $m$ agents with different strategies (e.g., different discount factors). In training mode, they individually interact with the real environment, generating execution latency data pool. These data are then integrated into a new execution buffer and employed to train the $\theta_{adv}$. Subsequently, they interact with the simulated environment to generate simulated experiences and concurrently update themselves. In inference mode, for a given query, $m$ agents independently generate candidate plans, utilize $\theta_{adv}$ to select the best plan, and the ultimately optimal plan is chosen from the $m$ best plans using $\theta_{adv}$.
We deploy FOSS with 2 agents. It demonstrates better performance than FOSS with 1 agent (Table II) and requires fewer epochs for convergence (Fig. 8). This can be attributed to the strategy of multi-agents, which not only enriches the diversity of training data for $\theta_{adv}$ but also generates a more diverse set of high-quality candidate plans. However, due to our serialization configuration, it incurs more training time than FOSS with 1 agent. If computational resources are ample, we recommend deploying FOSS with multi-agents in parallel.
@@ VII Related Work
Learned Query Optimizer. In recent years, various learned query optimizers have been proposed. Some works [4, 5, 7, 6, 10, 8, 9] model the plan generation process as MDP. Their primary goal is to train a learned value network to guide the bottom-up generation of more latency-effective plans. However, this method incurs high training costs due to its inefficient plan search, limiting its practicality. Other works [12, 26, 13, 27] aim to steer the plan generation in traditional optimizer using hint sets. This method is more practical, but its coarse-grained hint optimization approach cannot perform fine-grained optimization of join order and physical operators, limiting the plan search space. FOSS strikes a balance between these two approaches, offering practicality while still exploring a significant portion of the optimization possibilities.
Recently, the works [28, 29] build upon the foundation laid by Bao[12], employing hint sets similar to Bao. They focus on making the selection of candidate hint set more intelligent rather than relying solely on expert knowledge. Their work contributes positively to the advancement of plan-steerer.
Reinforcement Learning. Reinforcement learning can be divided into model-free RL and model-based RL. The former have experienced rapid development in the past few years, with many classic algorithms [30, 31, 21, 32, 33] being proposed and demonstrating excellent performance in various fields [34, 35, 36, 37]. However, model-free RL with high sample complexity are hard to be directly applied in real-world tasks, where trial-and-errors can be highly costly [18]. As a result, some scholars are dedicated to building simulated environment models to interact more efficiently with agents. FOSS follows the Dyna-style [38] method, using the learned environment model to generate more simulated experience data for the agent. In addition to the mentioned approach, in many model-based RL scenarios, the simulated environment models are differentiable. This enables policy learning through differential planning and value gradient methods [18].
Reinforcement Learning From Human Feedback. Reinforcement learning from human feedback (RLHF) is a technique employed for training AI systems to align with human goals. RLHF has emerged as the primary method for fine-tuning SOTA large language models [39]. The standard methodology for RLHF used today was introduced by [40] in 2017. Its framework is similar to FOSS, requiring feedback from external comparisons of two samples to train a reward model. This reward model is then used to train traditional RL, which is policy optimization. The key difference is that RLHF receives feedback from humans, making sample labeling more challenging due to human subjectivity, whereas FOSS obtains feedback from DBMS’s real execution.
@@
We introduce FOSS, a novel DRL-based query optimization system. Starting from the original plans generated by traditional optimizer, the planner of FOSS experiments with the optimization of these plans through step-by-step modifications, which results in candidate plans. We design an asymmetric advantage model to serve as a selector, choosing the optimal plan from the candidate plans. Furthermore, to enhance training efficiency, we employ the asymmetric advantage model as a reward indicator and integrate it with a traditional optimizer to build a simulated environment. With the simulated environment, FOSS can enhance its optimization ability through self-bootstrapping.
To our knowledge, FOSS is the pioneering learned-based query optimizer that achieves better-performance plans through fine-grained modifications to the original plans and employs a learned model for accelerating the training process of DRL. We believe that FOSS opens up a new path for query optimization and deserves further exploration.

### v2

@@ VI-B3 Optimization Time
We compare the optimization time of various optimizers on the entire JOB benchmark, which is the time from the input of SQL to the generation of the execution plan. As depicted in Fig. 7, we represent the comparative results using box plots. In contrast to PostgreSQL, learned-based optimizers often require longer optimization time due to operations like model inference and plan encoding. However, the overhead is negligible compared to the time saved in execution. Compared to Bao, Balsa, and HybridQO, FOSS has the smallest 25th percentile, 50th percentile, and 75th percentile of optimization time. The longer optimization time of FOSS compared to Loger is because the time taken by the traditional optimizer to generate the original plans accounts for a significant proportion of the optimization time in FOSS, while Loger does not require this. However, despite Loger having an advantage in optimization time, FOSS still surpasses it in total workload runtime.
@@
VI-B4 Known Best Plan
In this subsection, we analyze the maximum optimization potential for each learned query optimizer. We conduct three runs for each method on the entire JOB workload. From all the results, we acquire the known best plan (i.e., the one with the lowest execution latency) on each query for each learned optimizer. The comparison results are depicted in Fig. 8. It’s sorted by the time savings ratios of the known best plans relative to the original plans.
Overall, FOSS comprehensively outperforms Bao, Loger, and HybridQO, surpassing the expert optimizer across a wider range of queries at every percentile. Due to the limited plan search space of Bao, it generates better-performance plans than the expert optimizer across only a minimal range of queries. Despite Balsa implementing more fine-grained optimization by considering additional table scan operators, FOSS can still exhibit similar performance to Balsa on the top 65 ranked queries. This suggests that FOSS, even when focusing solely on join order and join methods optimization, achieves significant optimization effects. It also can be observed that due to the lack of assurance from the original plan, Balsa generates plans with poorer performance than the expert optimizer on 40 queries. Specially, FOSS, Balsa, HybridQO, Loger, and Bao achieve at least a 25% time savings on 62, 61, 38, 31 and 29 queries, respectively. They each save at least 75% of the time on 29, 29, 23, 20 and 14 queries, respectively.
@@ VI-C Analysis of design choices
2-Maxsteps 8.02 182.77 0.596
3-Maxsteps (FOSS) 9.09 209.01 0.436
4-Maxsteps 10.37 235.14 0.487
5-Maxsteps 10.75 259.45 0.470
Off-Simulated 48.01 210.47 0.691
Off-Penalty 8.65 208.29 0.465
Off-Validation 6.26 209.82 0.653
2-Agents 12.45 280.65 0.420
@@ VI-C1 Determination of Maxsteps
Larger $maxsteps$ implies that FOSS can explore more diverse plans through the modifications of longer sequences to the original plan. However, it also makes the training of FOSS more challenging. We set $maxsteps$ to 2, 3, 4, and 5, respectively, and compare their performances. As described in Table II, with an increase in $maxsteps$, the optimization time grows due to the consideration of more plans in one episode. Concurrently, training time gradually increases as the plan search space expands, and the number of plans to be executed grows. As illustrated in Fig. 9, when $maxsteps$ is set to 3, FOSS achieves the best performance with the minimum GMRL.
For a more in-depth comparison, as described in Fig. 7, we analyze the distribution of step status for known best plan, where step0 plans refers to the original plans, and stepN plans refers to plans that take $N$ steps from the original plans. In the bar for $2-maxsteps$, there is a significant number of step2 plans. It suggests that $2-maxsteps$ may be insufficient. In the bar for $5-maxsteps$, the proportion of step4 and step5 plans is relatively small. As demonstrated by the other two results, FOSS always yields effective plans within 1-3 steps.
The above results can be attributed to the fact that as $maxsteps$ increases, the longer action sequences complicate the agent’s exploration and make it more challenging for the AAM to select a good plan from the candidate plans. Therefore, considering both training complexity and experimental effectiveness, we set the $maxsteps$ to 3.
@@ VI-C2 Effect of Simulated Environment
To illustrate the effect of the simulated environment, we disable it in FOSS. It allows the agent to interact only with the real environment. Due to the extensive queries executed in the real environment, it significantly prolong the training time. To make training more feasible, we have to reduce the episode number for each agent update to 200. After 48 hours of training, despite benefiting from more reliable rewards feedback, its performance remains unsatisfactory. Its final GMRL is 0.691, which is significantly poorer than FOSS. The inadequacy can be ascribed to the insufficient exploration. It indicates the effectiveness of the simulated environment.
VI-C3 Effect of Penalty
Penalty measure in the reward function can guide FOSS to focus on generating valuable candidate plans. We set penalty coefficient to 0 in the default configuration. Table II illustrates that FOSS without penalty, although having a shorter training time compared to FOSS with penalty, performs worse. This can be attributed to FOSS with penalty allowing the planner to generate a more diverse set of candidate plans within a limited number of steps. It also enriches the training data for updating the AAM. This indicates that penalty measure has a positive impact on FOSS.
VI-C4 Effect of Promising Plans Validation
During planner training, FOSS collects the promising plans based on the feedback from the AAM and execute these plans when compute resources are idle. Then FOSS pushes their execution results into the execution buffer for updating the AAM. In this experiment, we remove this operation, meaning FOSS can only obtain training data by randomly sampling queries and collecting candidate plans. As described in Table II, it reduces training time but diminishes the performance. As shown in Fig. 9, its GMRL decreases slowly with training progress. The absence of validation for promising plans not only reduces diversity in the AAM training data, but also leads to the accumulation of estimation errors during planner training, as errors in the AAM are not corrected in a timely manner.
VI-C5 Training with Muti-Agents
To enhance the robustness of FOSS, we also incorporate a switch for multi-agents. FOSS can initialize $m$ agents with different strategies (e.g., different discount factors and learning rates). In training mode, they interact with the real environment, obtaining execution result. These data are then integrated into the execution buffer and employed to train the AAM. Subsequently, they individually interact with the simulated environment to generate simulated experiences and concurrently update themselves. In inference mode, for a given query, $m$ agents independently generate candidate plans, utilize AAM to select the best plan, and the ultimately execution plan is chosen from the $m$ best plans using AAM.
We deploy FOSS with 2 agents. It demonstrates better performance than FOSS with 1 agent (Table II) and requires fewer epochs for convergence (Fig. 9). This can be attributed to the strategy of multi-agents, which not only enriches the diversity of training data for AAM but also generates a more diverse set of high-quality candidate plans. However, due to our serialization configuration, it incurs more training time than FOSS with 1 agent. If computational resources are ample, we recommend deploying FOSS with multi-agents in parallel.
@@ VII Related Work
Learned Query Optimizer. In recent years, various learned query optimizers have been proposed. Apart from the two types of methods introduced in I, there are also other works aiming to replace the traditional cost model and incorporate appropriate plan search space reduction methods. For instance, Robopt [30] encodes plans to predict plan performance for cross-platform systems, basing the entire plan enumeration on a set of algebraic operations that operate on vectors. Leon [31] retains expert knowledge as much as possible. It follows the plan generation process of the expert optimizer and only replaces the traditional cost model with a learned calibration model initialized by the traditional cost model. Additionally, Leon employs another learned model to help reduce unnecessary plan space. These two learned models effectively balance immediate and long-term benefits for plan generation. Similarly, QPSeeker [32] also aims to replace the traditional cost model with a powerful learned model that learns from both data distribution and queries with the assistance of a language model. Additionally, it utilizes Monte Carlo Tree Search to provide an execution plan for the query.
There are also some works [33, 34] build upon the foundation laid by Bao[13], employing hint sets similar to Bao. They focus on making the selection of candidate hint set more intelligent rather than relying solely on expert knowledge. Their work contributes positively to the advancement of plan-steerer.
Reinforcement Learning. Reinforcement learning can be divided into model-free RL and model-based RL. The former have experienced rapid development in the past few years, with many classic algorithms [35, 36, 24, 37, 38] being proposed and demonstrating excellent performance in various fields [39, 40, 41, 42]. However, model-free RL with high sample complexity are hard to be directly applied in real-world tasks, where trial-and-errors can be highly costly [21]. As a result, some scholars are dedicated to building simulated environment to interact more efficiently with agents. FOSS follows the Dyna-style [43] method, using the learned environment model to generate more simulated experiences for the agent. In addition to the mentioned approach, in many model-based RL scenarios, the simulated environment models are differentiable. This enables policy learning through differential planning and value gradient methods [21].
@@ VIII Conclusion
We introduce FOSS, a novel DRL-based query optimization system. Starting from the original plans generated by traditional optimizer, the planner of FOSS experiments with the optimization of these plans through step-by-step modifications, which results in candidate plans. We design an asymmetric advantage model to serve as a selector, choosing the optimal one from the candidate plans. Furthermore, to enhance training efficiency, we employ the asymmetric advantage model as a reward indicator and integrate it with a traditional optimizer to build a simulated environment. With the simulated environment, FOSS can enhance its optimization ability through self-bootstrapping.
To our knowledge, FOSS is the pioneering learned-based query optimizer that achieves better-performance plans through fine-grained modifications to the original plans and efficiently self-bootstraps by interacting with the simulated environment model. We believe that FOSS opens up a new path for query optimization and deserves further exploration.
Acknowledgment
National Key Research & Develop Plan(2023YFB4503600); National Natural Science Foundation of China (U23A20299, 62072460, 62172424, 62276270, 62322214).